OpenStudy (anonymous):

write the equation of the line that has undefined slope and passes through the point(-5,9)

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

undefined means 0

OpenStudy (anonymous):

y-9=m(x+5) m= slope

OpenStudy (anonymous):

is undefined means 0

OpenStudy (anonymous):

no it means infinite

OpenStudy (anonymous):

so in place of m i have to put the sign of infinity

OpenStudy (anonymous):

here m=1/0 so for equation to exist x+5=0 wehave then y-9=0 that gives y=9

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

answer is x = -5

OpenStudy (anonymous):

slope undefined means the line is vertical. if it is vertical it looks like x = constant. since in this case x =-5 that is the answer.

OpenStudy (anonymous):

by the way, undefined does not mean infinite!

OpenStudy (anonymous):

it just means the line is vertical

OpenStudy (anonymous):

oops, yes satellite is right.